how many moles of sulfur dioxide are held inside a 67.4 l cansiter at a pressure of 0.25 atm and a temperature of 40.0C


1
Expert's answer
2021-05-24T07:21:31-0400

PV=nRT. n=PV/RT

R=0.0821 atm x L

Mol x K

T={273+40}K=310

P=0.25atm

V=67.4L


"n= \frac{(0.25atm)× (67.4L)}{0.0821atm × (310k)}=0.662mol"